# ConsRec: Denoising Sequential Recommendation through User-Consistent Preference Modeling

### v1

@@ 4 Experimental Methodology
Datasets. We utilize four sequential recommendation datasets–Beauty, Sports, Toys McAuley et al. (2015), and Yelp–to train and evaluate both the sequential recommendation module and the item filtering module. More details are shown in Appendix A.2.
@@
Evaluation Metrics. We adopt Recall@10/20 and NDCG@10/20 to evaluate the recommendation performance, which is consistent with prior works Xie et al. (2022); Liu et al. (2023); Chen et al. (2024). Statistical significance is determined using a permutation test with P$<0.05$.
@@
Baselines. We compare ConsRec with several widely used methods for sequential recommendation, including vanilla sequential recommendation models and some item denoising techniques.
Vanilla Sequential Recommendation Models. We compare ConsRec with both ID-based and content-based recommendation models.
@@
First, ConsRec is evaluated against seven ID-based recommendation models. GRU4Rec Hidasi et al. (2016) employs Recurrent Neural Networks (RNNs) to model user-item interaction sequences, while SASRec Kang and McAuley (2018) and Bert4Rec Sun et al. (2019) leverage self-attention to capture user preferences from sequential interactions. Additionally, we examine models that incorporate item-side information. S3Rec Zhou et al. (2020) enhances self-attention modules through four self-supervised pretraining strategies to better model relationships among attributes, items, and users. ICAI-SR Yuan et al. (2021b) represents item-attribute relationships using a heterogeneous graph to improve relevance modeling, whereas NOVA Liu et al. (2021a) integrates item attributes as side information via attention modules. DIF-SR Xie et al. (2022) introduces a non-invasive attention mechanism to fuse item attribute information into user behavior modeling effectively.
@@
Furthermore, ConsRec is compared with four content-based recommendation models. CoWPiRec Yang et al. (2023a) optimizes Pretrained Language Models (PLMs) to enhance item representation by aligning it with collaborative filtering representations derived from a word graph. TASTE Liu et al. (2023) leverages T5 Raffel et al. (2020) to directly encode the text representations of both users and items for matching items with user intentions. DWSRec Zhang et al. (2024b) leverages both fully whitened and relaxed whitened item representations as dual views for effective recommendations. Lastly, HLLM Chen et al. (2024) employs Large Language Models (LLMs) to generate item representations, subsequently utilizing a hierarchical architecture to predict the next item.
Item Denoising for Sequential Recommendation. We also evaluate ConsRec against three baselines that either explicitly remove noisy items or implicitly address item noise through denoising modeling. All these models are implemented based on the item-ID base recommendation model, such as Bert4Rec Sun et al. (2019). For methods that explicitly remove noisy items, DSAN Yuan et al. (2021a) learns a target item embedding and applies an adaptively sparse transformation function to filter out noisy interactions. Similarly, HSD Zhang et al. (2022a) identifies and excludes noisy interactions by detecting inconsistencies with user interests, leveraging the sequential order of recorded interactions. In contrast, DCRec Yang et al. (2023b) employs a debiased contrastive learning framework to implicitly mitigate the impact of noise.
@@
GRU4Rec Bert4Rec SASRec S3Rec NOVA ICAI-SR DIF-SR CoWPiRec DWSRec HLLM TASTE ConsRec
Beauty R@10 0.0530 0.0529 0.0828 0.0868 0.0887 0.0879 0.0908 0.0837 0.0989 0.0887 0.1030 $\text{}^{\dagger\ddagger}$
R@20 0.0839 0.0815 0.1197 0.1236 0.1237 0.1231 0.1284 0.1280 0.1496 0.1238 0.1550 $\text{}^{\dagger\ddagger}$
N@10 0.0266 0.0237 0.0371 0.0439 0.0439 0.0439 0.0446 0.0405 0.0470 0.0499 0.0517 $\text{}^{\dagger\ddagger}$
N@20 0.0344 0.0309 0.0464 0.0531 0.0527 0.0528 0.0541 0.0517 0.0597 0.0588 0.0649 $\text{}^{\dagger\ddagger}$
Sports R@10 0.0312 0.0295 0.0526 0.0517 0.0534 0.0527 0.0556 0.0588 0.0629 0.0555 0.0633 $\text{}^{\dagger\ddagger}$
R@20 0.0482 0.0465 0.0773 0.0758 0.0759 0.0762 0.0800 0.0892 0.0958 0.0816 0.0964 $\text{}^{\dagger\ddagger}$
N@10 0.0157 0.0130 0.0233 0.0249 0.0250 0.0243 0.0264 0.0300 0.0313 0.0295 0.0338 $\text{}^{\dagger\ddagger}$
N@20 0.0200 0.0173 0.0295 0.0310 0.0307 0.0302 0.0325 0.0376 0.0396 0.0360 0.0421 $\text{}^{\dagger\ddagger}$
Toys R@10 0.0370 0.0533 0.0831 0.0967 0.0978 0.0972 0.1013 0.0502 0.0967 0.0928 0.1232 $\text{}^{\dagger\ddagger}$
R@20 0.0588 0.0787 0.1168 0.1349 0.1322 0.1303 0.1382 0.0739 0.1472 0.1306 0.1789 $\text{}^{\dagger\ddagger}$
N@10 0.0184 0.0234 0.0375 0.0475 0.0480 0.0478 0.0504 0.0272 0.0457 0.0520 0.0640 $\text{}^{\dagger}$
N@20 0.0239 0.0297 0.0460 0.0571 0.0567 0.0561 0.0597 0.0331 0.0585 0.0616 0.0780 $\text{}^{\dagger}$
Yelp R@10 0.0361 0.0524 0.0650 0.0589 0.0681 0.0663 0.0698 0.0657 0.0702 0.0659 0.0738 $\text{}^{\dagger\ddagger}$
R@20 0.0592 0.0756 0.0928 0.0902 0.0964 0.0940 0.1003 0.0959 0.1103 0.0955 0.1156 $\text{}^{\dagger\ddagger}$
N@10 0.0184 0.0327 0.0401 0.0338 0.0412 0.0400 0.0419 0.0416 0.0382 0.0407 0.0397 $\text{}^{\ddagger}$
N@20 0.0243 0.0385 0.0471 0.0416 0.0483 0.0470 0.0496 0.0497 0.0482 0.0504 0.0502 $\text{}^{\dagger\ddagger}$
@@
Implementation Details. Our ConsRec is implemented using OpenMatch Liu et al. (2021b); Yu et al. (2023) and optimized using in-batch negatives. Both the item representation model in the filter module and the recommendation module are initialized from the T5-base checkpoint [^1] https://huggingface.co/google-t5/t5-base in Huggingface Transformers Wolf et al. (2019). The item representation model in the filter module is trained using the Adam optimizer with a learning rate of 1e-4 and a batch size of 32. For the recommendation module, we follow the settings in Liu et al. (2023), which is trained using the Adam optimizer with a learning rate of 1e-4, a warm-up proportion of 0.1, and a batch size of 8. We set the filtering threshold $\tau$ to 0.7, based on the hyperparameter sensitivity analysis reported in Appendix A.4. During inference, we excluded the last interaction from the user-interaction sequence when constructing the consistency graph, thereby avoiding potential data leakage.
5 Experimental Results
@@
In this section, we begin by evaluating the recommendation performance of ConsRec and perform ablation studies to assess the effectiveness of denoising modules and various training strategies. Then, we analyze text similarities among retained items to evaluate the effectiveness of the item representation learning method. Finally, case studies are shown in Appendix A.6.
5.1 Overall Performance
@@
In this section, we first implement ConsRec based on TASTE Liu et al. (2023) to evaluate its recommendation performance. Subsequently, we implement ConsRec with Bert4Rec Sun et al. (2019) to evaluate its denoising capability by comparing it against various baselines.
@@
Recommendation Performance. As shown in Table 1, we present the overall performance of ConsRec and various recommendation baseline models. Overall, ConsRec demonstrates its effectiveness by significantly outperforming all baselines across four datasets, leveraging TASTE Liu et al. (2023) as its backbone architecture.
@@
Among all baseline models, text-based recommendation models generally exhibit significantly stronger performance compared to item ID-based models. This superiority derives from the ability of PLMs and LLMs to effectively capture matching signals through text representations of users and items. Based on the TASTE backbone, ConsRec integrates a denoising module to remove noisy items, utilizing these retained user-item interactions to train the recommendation module. The evaluation results reveal a significant improvement in the recommendation performance, highlighting its effectiveness in enabling the recommendation module to better model user behavior. Among various recommendation tasks, ConsRec demonstrates much better performance on the Beauty and Sports datasets, while showing relatively weaker improvements on the Yelp dataset. This discrepancy may be attributed to the tendency of users to make more casual clicks driven by instant interest in these shopping scenarios.
@@
ConsRec 0.0398 0.0576 0.0211 0.0274
Table 2: Denoising Performance of Different Methods. To ensure a fair comparison, all denoising methods are implemented using Bert4Rec Sun et al. (2019).
Denoising Effectiveness. We compare ConsRec with three state-of-the-art denoising methods–DSAN Yuan et al. (2021a), HSD Zhang et al. (2022a), and DCRec Yang et al. (2023b)–to evaluate its denoising effectiveness. The baselines are categorized into two groups–explicit denoising methods (DSAN and HSD) and implicit denoising methods (DCRec). All methods are implemented using the item ID-based recommendation model, Bert4Rec Sun et al. (2019). The evaluation results are shown in Table 2.
Among the baseline methods, DCRec achieves the best performance by modeling user behaviors based on raw user-item interactions and effectively mitigating noise through dynamically assigning weights to user-interacted items. However, this implicit denoising approach necessitates model-specific training, restricting our experiments to the Bert4Rec model used in constructing the DCRec framework. Benefiting from explicit denoising methods, ConsRec can be applied to the item ID-based recommendation model and achieves a 4.7% improvement over all denoising baseline models, highlighting its effectiveness in filtering noisy items. This consistent improvement with the item ID-based model underscores the potential of extending the filtering approach of ConsRec to different architectures of recommendation models.
@@
Beauty ConsRec (Zero-Shot) 0.1631 0.0697
@@
w/o $M_{\text{Rec}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$
Sports ConsRec (Zero-Shot) 0.1014 0.0420
@@
w/o $M_{\text{Rec}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$
Toys ConsRec (Zero-Shot) 0.1797 0.0786
@@
w/o $M_{\text{Rec}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger\ddagger}$ $\text{}^{\dagger\ddagger}$
Yelp ConsRec (Zero-Shot) 0.1158 0.0504
@@
w/o $M_{\text{Filter}}$ Training $\text{}^{\dagger}$ $\text{}^{\dagger}$
@@
5.2 Ablation Study
@@
This section presents an ablation study to evaluate the contribution of the denoising module ($M_{\text{Filter}}$) and the effectiveness of our proposed training strategy in ConsRec. More experimental results are shown in Appendix A.5.
@@
Table 3 shows the performance of several ablated versions of our model across four datasets. Comparing ConsRec (w/o $M_{\text{Filter}}$) with ConsRec (Zero-Shot), we observe a consistent improvement in recommendation performance across all datasets when a zero-shot $M_{\text{Filter}}$ is applied during inference. This highlights the inherent benefit of even an untrained denoising module in refining user-interacted items for recommendation. The impact of the denoising module becomes more evident when comparing ConsRec (Zero-Shot) with ConsRec (w/o $M_{\text{Rec}}$ Training). Training the $M_{\text{Filter}}$ module using our instruction-based method leads to a significant improvement, demonstrating the effectiveness of learning tailored item representations through denoising. Specifically, ConsRec (w/o $M_{\text{Rec}}$ Training) consistently outperforms ConsRec (w/o $M_{\text{Filter}}$), indicating that a well-trained filter is crucial for recommendation systems.
@@
Furthermore, the full ConsRec model, which trains the recommendation module $M_{\text{Rec}}$ on the retained interactions, achieves the best performance across all datasets, with an average improvement of over 12% compared to ConsRec (Zero-Shot). This substantial gain underscores the importance of aligning the recommendation module with the retained data, enabling it to better capture user preferences and behaviors from the refined interaction history. The consistent statistically significant improvements of ConsRec over both ConsRec (w/o $M_{\text{Filter}}$) and ConsRec (Zero-Shot) further validate the effectiveness of our complete model.

### v2

@@ 4. Experimental Methodology
Datasets. We utilize four sequential recommendation datasets–Beauty, Sports, and Toys from Amazon Product (mcauley2015image), and Yelp–to train and evaluate both the sequential recommendation module and the semantic shortcut extractor. More details are shown in Appendix A.2.
@@
Evaluation Metrics. We adopt Recall@10/20 and NDCG@10/20 to evaluate the recommendation performance, which is consistent with prior works (Xie2022DIF-SR; liu2023text; chen2024hllm). Statistical significance is determined using a permutation test with $P<0.05$.
@@
Baselines. We compare LISRec with several widely used methods for sequential recommendation, including vanilla sequential recommendation models and some item denoising techniques.
@@
Vanilla Sequential Recommendation Models. We compare LISRec with both ID-based and content-based recommendation models. First, LISRec is evaluated against seven ID-based recommendation models. GRU4Rec (Hidasi2015session) employs Recurrent Neural Networks (RNNs) to model user-item interaction sequences, while SASRec (Wang2018self) and Bert4Rec (Sun2019Bert4rec) leverage self-attention to capture user preferences from sequential interactions. Additionally, we examine models that incorporate item-side information. S3Rec (Zhou2020s3) enhances self-attention modules through four self-supervised pretraining strategies to better model relationships among attributes, items, and users. ICAI-SR (yuan2021icai) represents item-attribute relationships using a heterogeneous graph to improve relevance modeling, whereas NOVA (Liu2021nova) integrates item attributes as side information via attention modules. DIF-SR (Xie2022DIF-SR) introduces a non-invasive attention mechanism to fuse item attribute information into user behavior modeling effectively.
@@
Furthermore, LISRec is compared with four content-based recommendation models. CoWPiRec (yang2023collaborative) optimizes Pretrained Language Models (PLMs) to enhance item representation by aligning it with collaborative filtering representations derived from a word graph. TASTE (liu2023text) leverages T5 (raffel2020exploring) to directly encode the text representations of both users and items for matching items with user intentions. DWSRec (zhang2024dual) leverages both fully whitened and relaxed whitened item representations as dual views for effective recommendations. Lastly, HLLM (chen2024hllm) employs Large Language Models (LLMs) to generate item representations, subsequently utilizing a hierarchical architecture to predict the next item.
Table 1. Overall Performance of LISRec. The best results are marked in bold, while the second-best results are underlined. The HLLM model is reproduced using a 1B-scale LLM. $\dagger$} and $\ddagger$} denote statistically significant improvements over $\text{DIF-SR}^{\dagger}$ and $\text{TASTE}^{\ddagger}$, respectively.
@@
GRU4Rec Bert4Rec SASRec S3Rec NOVA ICAI-SR DIF-SR CoWPiRec DWSRec HLLM TASTE LISRec
@@
[…]
Item Denoising for Sequential Recommendation. We also evaluate LISRec against three baselines that either explicitly remove noisy items or implicitly address item noise through denoising modeling. All these models are implemented based on the item-ID base recommendation model, such as Bert4Rec (Sun2019Bert4rec). For methods that explicitly remove noisy items, DSAN (yuan2021dual) learns a target item embedding and applies an adaptively sparse transformation function to filter out noisy interactions. Similarly, HSD (zhang2022hierarchical) identifies and excludes noisy interactions by detecting inconsistencies with user interests, leveraging the sequential order of recorded interactions. In contrast, DCRec (yang2023debiased) employs a debiased contrastive learning framework to implicitly mitigate the impact of noise.
@@
Implementation Details. LISRec is implemented using OpenMatch [^1] https://github.com/OpenMatch/OpenMatch (liu2021openmatch; yu2023openmatch) and optimized using in-batch negatives. Both the item representation model in the semantic shortcut extractor and the recommendation module are initialized from the T5-base checkpoint [^2] https://huggingface.co/google-t5/t5-base in Huggingface Transformers (wolf2019huggingface). The item representation model in the semantic shortcut extractor is trained using the Adam optimizer with a learning rate of 1e-4 and a batch size of 32. For the recommendation module, we follow the settings in liu2023text, which is trained using the Adam optimizer with a learning rate of 1e-4, a warm-up proportion of 0.1, and a batch size of 8. We set the filtering threshold $\tau$ to 0.7, based on the hyperparameter sensitivity analysis reported in Appendix 5.3. During inference, we excluded the last interaction from the user-interaction sequence when constructing the consistency graph, thereby avoiding potential data leakage.
Additionally, we reproduce all item-ID based models and DWSRec (zhang2024dual) using implementations provided by RecBole [^3] https://recbole.io/ (zhao2021recbole), maintaining consistent settings across all models for fair comparison. For CoWPiRec and TASTE, we adhere to the default experimental settings as specified in their original papers. For HLLM, due to computational constraints, we ultilize TinyLlama-1.1B checkpoint [^4] https://huggingface.co/TinyLlama/TinyLlama-1.1B-intermediate-step-1431k-3T (zhang2024tinyllama) as the item-LLM component, and the user-LLM is implemented as a custom Llama model (touvron2023llama) with two hidden layers, trained from scratch using the following settings: learning rate of 1e-4, batch size of 2, training for 5 epochs, and a negative sampling ratio of 1024 negative samples per positive sample.
5. Experimental Results
@@
In this section, we begin by evaluating the recommendation performance of LISRec and perform ablation studies to assess the effectiveness of the semantic shortcut extractor and various training strategies. Then, we analyze textual similarities among retained items to evaluate the effectiveness of the item representation learning method. Finally, several case studies are shown.
5.1. Overall Performance
@@
In this section, we first implement LISRec based on TASTE (liu2023text) to evaluate its recommendation performance. Subsequently, we implement LISRec with Bert4Rec (Sun2019Bert4rec) to evaluate its denoising capability by comparing it against various baselines.
5.1.1. Recommendation Performance
@@
As shown in Table 1, we present the overall performance of LISRec and various recommendation baseline models. Overall, LISRec demonstrates its effectiveness by significantly outperforming all baselines across four datasets, leveraging TASTE (liu2023text) as its backbone architecture.
@@
Among all baseline models, text-based recommendation models generally exhibit significantly stronger performance compared to item ID-based models. This superiority derives from the ability of PLMs and LLMs to effectively capture matching signals through text representations of users and items. Based on the TASTE backbone, LISRec integrates a semantic shortcut extractor to remove noisy items, utilizing these retained user-item interactions to train the recommendation module. The evaluation results reveal a significant improvement in the recommendation performance, highlighting its effectiveness in enabling the recommendation module to better model user behavior. Among various recommendation tasks, LISRec demonstrates much better performance on the Beauty and Sports datasets, while showing relatively weaker improvements on the Yelp dataset. This discrepancy may be attributed to the tendency of users to make more casual clicks driven by instant interest in these shopping scenarios.
Table 2. Denoising Performance of LISRec. To ensure a fair comparison, all denoising methods are implemented using Bert4Rec (Sun2019Bert4rec) as the backbone model.
@@
LISRec 0.0398 0.0576 0.0211 0.0274
5.1.2. Denoising Effectiveness
We compare LISRec with three state-of-the-art denoising methods–DSAN (yuan2021dual), HSD (zhang2022hierarchical), and DCRec (yang2023debiased)–to evaluate its denoising effectiveness. The baselines are categorized into two groups–explicit denoising methods (DSAN and HSD) and implicit denoising methods (DCRec). All methods are implemented using the item ID-based recommendation model, Bert4Rec (Sun2019Bert4rec). The evaluation results are shown in Table 2.
Among the baseline methods, DCRec achieves the best performance by modeling user behaviors based on raw user-item interactions and effectively mitigating noise through dynamically assigning weights to user-interacted items. However, this implicit denoising approach necessitates model-specific training, restricting our experiments to the Bert4Rec model used in constructing the DCRec framework. Benefiting from explicit denoising methods, LISRec can be applied to the item ID-based recommendation model and achieves a 4.7% improvement over all denoising baseline models, highlighting its effectiveness in filtering noisy items. This consistent improvement with the item ID-based model underscores the potential of extending the filtering approach of LISRec to different architectures of recommendation models.
Table 3. Ablation Study of LISRec. $\dagger$} and $\ddagger$} denote statistically significant improvements over $\text{LISRec{} (w/o $M_{\text{Filter}}$)}^{\dagger}$ and $\text{LISRec{} (Zero-Shot)}^{\ddagger}$.
@@ 5.1.2. Denoising Effectiveness
Beauty LISRec (Zero-Shot) 0.1631 0.0697
@@
[…]
Sports LISRec (Zero-Shot) 0.1014 0.0420
@@
[…]
Toys LISRec (Zero-Shot) 0.1797 0.0786
@@
[…]
Yelp LISRec (Zero-Shot) 0.1158 0.0504
@@
Table 4. Ablation Study on Different Filter Models in LISRec.
Dataset Method R@20 N@20
Beauty LISRec 0.1774 0.0772
w/ Vanilla T5 $M_{\text{Filter}}$ 0.1761 0.0762
w/ TASTE $M_{\text{Filter}}$ 0.1767 0.0767
Sports LISRec 0.1082 0.0496
w/ Vanilla T5 $M_{\text{Filter}}$ 0.1073 0.0457
w/ TASTE $M_{\text{Filter}}$ 0.1077 0.0471
Toys LISRec 0.1896 0.0826
w/ Vanilla T5 $M_{\text{Filter}}$ 0.1879 0.0802
w/ TASTE $M_{\text{Filter}}$ 0.1883 0.0811
Yelp LISRec 0.1173 0.0509
w/ Vanilla T5 $M_{\text{Filter}}$ 0.1169 0.0507
w/ TASTE $M_{\text{Filter}}$ 0.1170 0.0508
5.2. Ablation Study
@@
This section presents an ablation study to evaluate the contribution of the semantic shortcut extractor ($M_{\text{Filter}}$) and the effectiveness of our proposed training strategy in LISRec.
@@
Table 3 shows the performance of several ablated versions of our model across four datasets. Comparing LISRec (w/o $M_{\text{Filter}}$) with LISRec (Zero-Shot), we observe a consistent improvement in recommendation performance across all datasets when a zero-shot $M_{\text{Filter}}$ is applied during inference. This highlights the inherent benefit of even an untrained semantic shortcut extractor in refining user-interacted items for recommendation. The impact of the semantic shortcut extractor becomes more evident when comparing LISRec (Zero-Shot) with LISRec (w/o $M_{\text{Rec}}$ Training). Training the $M_{\text{Filter}}$ module with our instruction-based approach results in significant performance gains, demonstrating the value of learning tailored item representations for extracting semantic shortcuts. In particular, LISRec (w/o $M_{\text{Rec}}$ Training) consistently outperforms LISRec (w/o $M_{\text{Filter}}$), highlighting that an effective semantic shortcut extractor is essential for improved recommendation accuracy.
@@
Furthermore, the full LISRec model, which trains the recommendation module $M_{\text{Rec}}$ on the retained interactions, achieves the best performance across all datasets, with an average improvement of over 12% compared to LISRec (Zero-Shot). This substantial gain underscores the importance of aligning the recommendation module with the retained data, enabling it to better capture user preferences and behaviors from the refined interaction history. The consistent statistically significant improvements of LISRec over both LISRec (w/o $M_{\text{Filter}}$) and LISRec (Zero-Shot) further validate the effectiveness of our complete model.
To evaluate the effectiveness of the $M_{\text{Filter}}$ module in LISRec, we compare it with two alternatives: a vanilla T5 $M_{\text{Filter}}$ initialized from T5-base without task-specific training, and a TASTE $M_{\text{Filter}}$ (trained on recommendation data).
As shown in Table 4, LISRec achieves the best performance across all datasets, consistently surpassing both baselines. The vanilla T5 $M_{\text{Filter}}$ lags behind due to its lack of adaptation to recommendation data, while the TASTE-based variant, though improved, is still limited by its exposure to only a single dataset during training. In contrast, our approach benefits from balanced pre-training over multiple datasets, leading to more robust and transferable item representations. These results highlight the importance of both domain-specific and multi-dataset training for semantic shortcut extraction in recommendation.
(a) Bert4Rec
(b) TASTE
Figure 3. Hyperparameter Sensitivity Analysis. We conduct experiments on the threshold hyperparameter, $\tau$, with results derived from the Amazon Beauty dataset using the Bert4Rec 3(a) and TASTE 3(b) backbone models.
5.3. Hyperparameter Sensitivity Analysis
We analyze the impact of the similarity threshold $\tau$, which is a key hyperparameter in the construction of the personal item graph described in Sec.3.2.2, by varying $\tau$ from 0.2 to 0.8 while keeping all other experimental settings fixed. Figure 3 shows the performance of Bert4Rec and TASTE across four evaluation metrics (Recall@10/20 and NDCG@10/20).
The results reveal a clear and consistent pattern for both backbone models. When $\tau$ is set too low (0.2 to 0.4), the semantic shortcut extractor becomes overly permissive, allowing many irrelevant or noisy items into the shortcut, which results in degraded performance. As $\tau$ increases to a moderate range (0.5 to 0.6), performance improves markedly, indicating that the shortcut extractor is now able to focus on truly relevant interactions and effectively filter out noise. For example, Bert4Rec achieves its best R@10 around $\tau=0.6$, and similar patterns are observed for all metrics and for the TASTE backbone.
Importantly, our LISRec framework consistently outperforms baseline models without semantic filtering across the entire range of $\tau$ values. This demonstrates the robustness and generalizability of our approach–even when the threshold is not optimally tuned, shortcut extraction leads to reliable improvements in recommendation quality.
(a) Raw User-Item Interaction Trained $M_{\text{Rec}}$.
(b) $M_{\text{Filter}}$-Retained Items Trained $M_{\text{Rec}}$.
Figure 4. Performance of Recommendation Models Using Different Semantic Shortcut Extraction Methods.
5.4. Effectiveness of Different Item Representations for Shortcut Extraction
To evaluate the effectiveness of the semantic shortcut extractor in LISRec, we perform a comparative analysis involving various item representation strategies. Specifically, we compare the full LISRec model with its variant without the semantic shortcut extractor (LISRec w/o Filter), as well as versions of LISRec equipped with different item representation methods within the semantic shortcut extractor.
We implement the semantic shortcut extractor $M_{\text{Filter}}$ using three approaches: an ID-based recommendation model, Bert4Rec (Sun2019Bert4rec), and a dense retriever, BGE-Base (bge_embedding), which serve as baselines for evaluating representation effectiveness.